OpenStudy (anonymous):

a 90 kg cyclist riding a 10 kg bike accelerates from 0 m/s to 8 m/s in 3 seconds. what is his acceleration?

geerky42 (geerky42):

\(a = \dfrac{v_f-v_o}{t}\) So we have \(a = \dfrac{8~m/s - 0~m/s}{3s} = \boxed{\dfrac{8}{3}m/s^2}\)

geerky42 (geerky42):

Well, that's the acceleration of bike, but since he is sitting on bike, riding along with it, his acceleration is same.

OpenStudy (anonymous):

thank you (: